# Weighted Ehrhart theory via mixed Hodge modules on toric varieties

### v1

@@ 1.1. Combinatorial and geometrical posets and their identification
Let $M\cong{\mathbb{Z}}^{n}$ be a lattice and $P\subset M_{{\mathbb{R}}}:=M\otimes{\mathbb{R}}\cong{\mathbb{R}}^{d}$ be a full-dimensional lattice polytope. To $P$ one associates a fan $\Sigma=\Sigma_{P}\subset N_{\mathbb{R}}:=N\otimes{\mathbb{R}}$, with $N$ the dual lattice of $M$, called the inner normal fan of $P$. To the fan $\Sigma_{P}$ one further associates a projective toric variety $X=X_{P}$ with torus ${\mathbb{T}}$ (whose character lattice is $M$) and an ample Cartier divisor $D=D_{P}$. There is a order-reversing one-to-one correspondence between the faces $\emptyset\neq Q\preceq P$ of $P$ and the cones $\sigma_{Q}$ of $\Sigma_{P}$, with $\dim_{\mathbb{R}}(Q)=n-\dim_{\mathbb{R}}(\sigma_{Q})$. There is an order-reversing one-to-one correspondence between the cones $\sigma_{Q}$ of $\Sigma_{P}$ and the ${\mathbb{T}}$-orbits $O_{Q}$ of the toric variety $X_{P}$, with $\dim_{\mathbb{C}}(O_{Q})=n-\dim_{\mathbb{R}}(\sigma_{Q})$. Then the identification between the faces $\emptyset\neq Q\preceq P$ of $P$ and the torus orbits $O_{Q}$ of the toric variety $X_{P}$ is order-preserving, with $\dim_{\mathbb{R}}(Q)=\dim_{\mathbb{C}}(O_{Q})$.
@@ 1.3. Geometric viewpoint on weighted Ehrhart theory
Then the classical Dehn-Sommerville relations correspond to the Poincaré duality isomorphism for rational cohomology, motivating the use of intersection cohomology coefficients for non-simple lattice polytopes. In fact, using the intersection Hodge module ${IC^{\prime}}^{H}_{X_{P}}$ and its purity, we give a geometric proof of formula (1.4), as it will be explained below.
@@ 1.3.2. Degree calculations
As an example, ${\mathcal{M}}={IC^{\prime}}^{H}_{X}$ is pure of weight zero in the sense of Saito [22, 23], so that $IH^{i}(X):=H^{i}(X;{IC^{\prime}}^{H}_{X})$ is a pure Hodge structure of weight $i$. On the other hand, $[IH^{*}(X)]\in{\mathbb{Z}}[u^{\pm 1}]$, i.e., it is of Tate type. These two properties imply the vanishing of odd intersection cohomology groups, so that $I\chi_{y}(X):=\chi_{y}([IH^{*}(X)])$ coincides with the intersection cohomology Poincaré polynomial $IP_{X}(t)=h_{P}(t)$, up to the change of variable $t^{2}=-y$. As is well-known, one also has a similar purity and odd vanishing property for the stalks of $[{IC^{\prime}}^{H}_{X}]$ (e.g., see [25]). So we get
(1.7) $\begin{split}h_{P}(t)&=IP_{X}(t)=E_{P,f}(0,-t^{2})\\ &=\sum_{Q\preceq P}\widetilde{g}_{Q}(t^{2})\cdot(t^{2}-1)^{\dim(Q)}\\ &=\sum_{\emptyset\preceq Q^{\circ}\prec P^{\circ}}{g}_{Q^{\circ}}(t^{2})\cdot(t^{2}-1)^{n-1-\dim(Q^{\circ})},\end{split}$
@@
This already recovers Stanley’s “master duality” theorem [28, Thm. 3.14.9], as well as the generalized Dehn-Sommerville relations for the $h$-polynomial of a lattice polytope (boiling down to Poincaré duality for $IH^{*}(X)$). In terms of the corresponding Ehrhart polynomials, this duality formula reads as
@@ 2. Homology Hirzebruch classes via mixed Hodge modules
commuting with proper pushforwards, with $Gr^{F}_{p}\hbox{\rm DR}({\mathcal{M}})\simeq 0$ for almost all $p$ and ${\mathcal{M}}$ fixed; see [22, Sec.2.3], [23, Sec.3.10, Prop.3.11] and [24, Sec.1]. Here $D^{b}_{\rm coh}(X)$ denotes the bounded derived category of ${\mathcal{O}}_{X}$-modules with coherent cohomology.
@@
$\begin{split}\hbox{\rm DR}_{y}([M])&:=\sum_{i,p}\,(-1)^{i}\big{[}{\mathcal{H}}^{i}Gr^{F}_{-p}\hbox{\rm DR}(M)\big{]}\cdot(-y)^{p}\\ &=\sum_{p}\,\big{[}Gr^{F}_{-p}\hbox{\rm DR}(M)\big{]}\cdot(-y)^{p}\end{split}$
@@
(2.1) $\chi_{y}(H^{\bullet}):=DR_{y}([H^{\bullet}])=T_{y*}([H^{\bullet}])=\sum_{j,p}\,(-1)^{j}\dim_{{\mathbb{C}}}\text{\rm Gr}_{F}^{p}H_{{\mathbb{C}}}^{j}\,(-y)^{p}$
@@
If $X$ is smooth and $M={\mathbb{Q}}_{X}^{H}$ is the constant Hodge module, then
@@
This follows, as in [19], from [17, Proposition 4.2], via the identification of these Ishida sheaves with the (shifted) graded parts of the filtered Du Bois complex (and similarly for the Du Bois complex of a complex algebraic $V$-manifold, by [20, Section 2.7]). ∎
@@
As shown in [26, Corollary 5.19, Remark 5.20], the transformations $\hbox{\rm DR}_{y}$ and $T_{y*}$ commute with duality, i.e.,
@@
and similarly for $T_{y*}$. A similar statement holds for the motivic transformations $mC_{y}$ and $T_{y*}$, with duality on the localization $K_{0}(var/X)[{\mathbb{L}}^{-1}]$ defined as in [26, Section 4B] and [3]. All these duality transformations are functorial for proper maps, so that for $X$ compact with $k:X\to pt$ proper, one gets
@@
In the case of a projective complex algebraic variety $X$, one has by the intersection cohomology Hodge index theorem [21, Sect.3.6] the equality
@@ 3. Computations via stratifications
Under the assumption of (virtual) stratum-wise constancy, formula (3.6) reduces the calculation of Hodge-Chern and Hirzebruch classes to the motivic calculus of [19], up to assigning the above mentioned Laurent polynomial weights $\chi_{y}(i_{s}^{*}[{\mathcal{M}}])$ to each stratum $S$, with $s\in S$ chosen. For instance, the stratum-wise constancy assumption is automatically satisfied in the toric context for ${\mathbb{T}}$-equivariant mixed Hodge modules, e.g., the $IC$-complex, see [29, Lemma 1.2].
@@
More generally, the property of stratum-wise constant higher direct image sheaves $R^{\ell}f_{*}{\mathbb{Q}}_{Y}|_{S}$ is satisfied if $f:f^{-1}(S)\to S$ is a Zariski locally trivial fibration for all strata $S\in{\mathcal{S}}$. This property also holds for a ${\mathbb{T}}$-equivariant proper morphism $f:Y\to X$, with $X$ a toric variety with torus ${\mathbb{T}}$ (by [29, Lemma 1.2]).
The terms $mC_{y}([S\hookrightarrow{\bar{S}}])=\hbox{\rm DR}_{y}\big{[}(i_{S,{\bar{S}}})_{!}{\mathbb{Q}}_{S}^{H}\big{]}$ appearing in formula (3.6) are computed via resolutions of singularities as follows (see [20, Prop.5.2.1] and [19, Prop.2.2]):
@@
(3.8) $mC_{y}([S\hookrightarrow{\bar{S}}])=\hbox{\rm DR}_{y}(\big{[}(i_{S,{\bar{S}}})_{!}{\mathbb{Q}}^{H}_{S}\big{]})=(\pi_{Z})_{*}\big{[}\mathcal{O}_{X}(-D)\otimes\Lambda_{y}\Omega_{Z}^{1}(\log(D)\big{]}\in K_{0}({\bar{S}})[y],$
@@
For later use, let us also mention here the following formula, using the above notations (see, e.g., [26, Example 5.8]):
@@ 5. Characteristic class formulae in toric geometry
It suffices to prove the above properties on each open affine subvariety $U_{\sigma}\subset X$, for $\sigma\in\Sigma$, with the splitting $U_{\sigma}=Z_{\sigma}\times O_{\sigma},$ as recalled above. Let $p_{\sigma}:U_{\sigma}\to Z_{\sigma}$ be the smooth projection map with fiber $O_{\sigma}$. Then $(a)$ follows from the smooth base change property for the underlying constructible complex. Similarly, smooth pullback of mixed Hodge module complexes preserves the $K_{0}({\rm MHS}^{p})^{\rm Tate}$-module structure. Item $(b)$ can also be checked on the underlying perverse sheaf $IC_{X}$, where it follows by induction using $(a)$ and the Deligne construction. Again, these operations preserve the $K_{0}({\rm MHS}^{p})^{\rm Tate}$-module structure (see also [25, (1.7.6)]). For $(c)$, the stratum-wise constant case follows inductively from the fact that duality commutes (up to a shift) with the smooth pullback. Applying this to $[(j_{S})_{!}{\mathbb{Q}}^{H}_{X}]$ as in Example 3.5, one proves the virtual stratum-wise constant case since duality commutes with cross-products (and duality on a point preserves the Tate type). ∎
@@ 6. Applications to weighted lattice point counting
For a face $Q$ of $P$, denote by $i_{Q}:V_{\sigma_{Q}}:=X_{Q}\hookrightarrow X$ the inclusion of the orbit closure associated to the (cone $\sigma_{Q}$ of the) face $Q$. Note that we have $\dim_{\mathbb{R}}(Q)=\dim_{\mathbb{R}}(O_{\sigma_{Q}}).$ Then by (5.4) the following equality holds:
$\int_{X}ch({\mathcal{O}}_{X}(\ell D))\cap T_{y*}([{\mathcal{M}}])=\sum_{Q\preceq P}\chi_{y}(i_{x_{Q}}^{*}[{\mathcal{M}}])\cdot(1+y)^{\dim(Q)}\int_{X}ch({\mathcal{O}}_{X}(\ell D))\cap(i_{Q})_{*}td_{*}([\omega_{X_{Q}}]).$
@@
(6.7) $\int_{X}ch({\mathcal{O}}_{X}(\ell D))\cap(i_{Q})_{*}td_{*}([\omega_{X_{Q}}])=|{\rm Relint}(\ell Q)\cap M|.$
@@
(6.8) $\int_{X}ch({\mathcal{O}}_{X}(\ell D))\cap(i_{Q})_{*}td_{*}([\omega_{X_{Q}}])=\int_{X_{Q}}ch({\mathcal{O}}_{X_{Q}}(\ell D_{Q}))\cap td_{*}([\omega_{X_{Q}}]).$
The latter equality uses the projection formula, together with $(i_{Q})^{*}({\mathcal{O}}_{X_{P}}(\ell D_{P})={\mathcal{O}}_{X_{Q}}(\ell D_{Q})$. ∎
@@
Note that any ${\mathcal{M}}\in D^{b}{\rm MHM}(X_{P})$ (resp., $[{\mathcal{M}}]\in K_{0}({\rm MHM}(X_{P}))$), a (Grothendieck class of a) mixed Hodge module complex on $X_{P}$ which is (virtually) stratum-wise constant with respect to the stratification given by the torus orbits, induces a weight function on the faces of $P$ defined by $f_{Q}(y)=\chi_{y}(i_{x_{Q}}^{*}[{\mathcal{M}}])$, with $x_{Q}\in O_{\sigma_{Q}}$ a point in the orbit of $X_{P}$ associated to (the cone $\sigma_{Q}$ in the inner normal fan of $P$ corresponding to) the face $Q$ and $i_{x_{Q}}:\{x_{Q}\}\to X_{P}$ the inclusion map. Moreover, any weight function $f$ can be obtained in this way from such a mixed Hodge module complex ${\mathcal{M}}$, e.g., by using direct sums and shifts of $(i_{Q})_{!}{\mathbb{Q}}^{H}_{O_{\sigma_{Q}}}(n_{Q})$, with $n_{Q}\in{\mathbb{Z}}$ (or, using the subgroup $\chi_{Hdg}\left(\tau(F^{{\mathbb{T}}}(X)[y^{\pm 1}])\right)$ as in the Introduction). In particular, the expression $E_{P,f}(\ell,y)$ is obtained by evaluating a polynomial $E_{P,f}(z,y)$ at $z=\ell$.
@@
Let us next give some examples for Proposition 6.1, formulated in terms of Ehrhart theory. If $X=X_{P}$ is the projective toric variety with inner normal fan $\Sigma=\Sigma_{P}$ and ample Cartier divisor $D=D_{P}$ associated to a full-dimensional lattice polytope $P\subset M_{{\mathbb{R}}}\cong{\mathbb{R}}^{n}$, and $\ell\in{\mathbb{Z}}_{>0}$, for a face $Q$ of $P$, denote by $i_{Q}:V_{\sigma_{Q}}:=X_{Q}\hookrightarrow X$ the inclusion of the orbit closure associated to the (cone of the) face $Q$.
@@
    For ${\mathcal{M}}=(i_{Q})_{!}{\mathbb{Q}}^{H}_{O_{\sigma_{Q}}}$, we get
@@
    For ${\mathcal{M}}=(i_{Q})_{*}{\mathbb{Q}}^{H}_{O_{\sigma_{Q}}}$, we get by (5.2) and Danilov’s formula (6.1) that
@@
with $Q$ a face of $P$ and $Q^{\circ}$ the corresponding face of the polar polytope $P^{\circ}$, and $g_{Q^{\circ}}(t)\in{\mathbb{Z}}[t]$ the Stanley $g$-polynomial of $Q^{\circ}$, cf. [27, Sect.3]. We then get:
@@
(6.23) $\begin{split}\hbox{\rm DR}_{y}([{\mathcal{M}}])&=\sum_{Q\preceq P}\,\chi_{y}(i^{*}_{x_{Q}}[{\mathcal{M}}])\cdot(1+y)^{\dim(Q)}\cdot(i_{Q})_{*}[\omega_{V_{\sigma_{Q}}}]\\ &=\sum_{Q\preceq P}\,f_{Q}(y)\cdot\hbox{\rm DR}_{y}\left([(i_{Q})_{!}{\mathbb{Q}}^{H}_{O_{\sigma_{Q}}}]\right),\end{split}$
@@
$\hbox{\rm DR}_{y}\left({\mathbb{D}}_{X}\left([(i_{Q})_{!}{\mathbb{Q}}^{H}_{O_{\sigma_{Q}}}]\right)\right)=(-y)^{-\dim(Q)}\cdot\hbox{\rm DR}_{y}\left([(i_{Q})_{*}{\mathbb{Q}}^{H}_{O_{\sigma_{Q}}}]\right).$
@@
$\hbox{\rm DR}_{1/y}({\mathbb{D}}_{X}[{\mathcal{M}}])=\sum_{Q\preceq P}\,f_{Q}({y})\cdot(-y)^{\dim(Q)}\cdot\hbox{\rm DR}_{1/y}\left([(i_{Q})_{*}{\mathbb{Q}}^{H}_{O_{\sigma_{Q}}}]\right).$
@@
The reciprocity formula (6.20) can be applied to all cases of Example 6.5. In the (dual) cases $(a)$ and $(c)$, for ${\mathcal{M}}=(i_{Q})_{!}{\mathbb{Q}}^{H}_{O_{\sigma_{Q}}}$ with ${\mathbb{D}}_{X}[{\mathcal{M}}]=[(i_{Q})_{*}{\mathbb{Q}}^{H}_{O_{\sigma_{Q}}}(\dim(Q))]\in K_{0}({\rm MHM}(X))$, formula (6.20) specializes to the classical Ehrhart reciprocity (6.3) for $Q$. Indeed,

### v2

@@ 1.1. Combinatorial and geometrical posets and their identification
Let $M\cong{\mathbb{Z}}^{n}$ be a lattice and $P\subset M_{{\mathbb{R}}}:=M\otimes{\mathbb{R}}\cong{\mathbb{R}}^{n}$ be a full-dimensional lattice polytope. To $P$ one associates a fan $\Sigma=\Sigma_{P}\subset N_{\mathbb{R}}:=N\otimes{\mathbb{R}}$, with $N$ the dual lattice of $M$, called the inner normal fan of $P$. To the fan $\Sigma_{P}$ one further associates a projective toric variety $X=X_{P}$ with torus ${\mathbb{T}}$ (whose character lattice is $M$) and an ample Cartier divisor $D=D_{P}$. There is a order-reversing one-to-one correspondence between the faces $\emptyset\neq Q\preceq P$ of $P$ and the cones $\sigma_{Q}$ of $\Sigma_{P}$, with $\dim_{\mathbb{R}}(Q)=n-\dim_{\mathbb{R}}(\sigma_{Q})$. There is an order-reversing one-to-one correspondence between the cones $\sigma_{Q}$ of $\Sigma_{P}$ and the ${\mathbb{T}}$-orbits $O_{Q}$ of the toric variety $X_{P}$, with $\dim_{\mathbb{C}}(O_{Q})=n-\dim_{\mathbb{R}}(\sigma_{Q})$. Then the identification between the faces $\emptyset\neq Q\preceq P$ of $P$ and the torus orbits $O_{Q}$ of the toric variety $X_{P}$ is order-preserving, with $\dim_{\mathbb{R}}(Q)=\dim_{\mathbb{C}}(O_{Q})$.
@@ 1.3. Geometric viewpoint on weighted Ehrhart theory
Then the classical Dehn-Sommerville relations correspond to the Poincaré duality isomorphism for rational cohomology, motivating the use of intersection cohomology coefficients for non-simple lattice polytopes.
In this paper, we give a geometric interpretation and proofs of the above properties. For instance, using the intersection Hodge module ${IC^{\prime}}^{H}_{X_{P}}$ and its purity, we give a geometric proof of formula (1.4), as it will be explained below. For generalizations to the equivariant context, which will imply these properties for weighted Ehrhart polynomials depending on any homogeneous polynomial functions $\varphi:M_{\mathbb{R}}\to{\mathbb{C}}$ as in [2], see the authors’ paper [20].
@@ 1.3.2. Degree calculations
As an example, ${\mathcal{M}}={IC^{\prime}}^{H}_{X}$ is pure of weight zero in the sense of Saito [23, 24], so that $IH^{i}(X):=H^{i}(X;{IC^{\prime}}^{H}_{X})$ is a pure Hodge structure of weight $i$. On the other hand, $[IH^{*}(X)]\in{\mathbb{Z}}[u^{\pm 1}]$, i.e., it is of Tate type. These two properties imply the vanishing of odd intersection cohomology groups, so that $I\chi_{y}(X):=\chi_{y}([IH^{*}(X)])$ coincides with the intersection cohomology Poincaré polynomial $IP_{X}(t)=h_{P}(t^{2})$, up to the change of variable $t^{2}=-y$. As is well-known, one also has a similar purity and odd vanishing property for the stalks of $[{IC^{\prime}}^{H}_{X}]$ (e.g., see [26]). So we get
(1.7) $\begin{split}h_{P}(t^{2})&=E_{P,f}(0,-t^{2})\\ &=\sum_{Q\preceq P}\widetilde{g}_{Q}(t^{2})\cdot(t^{2}-1)^{\dim(Q)}\\ &=\sum_{\emptyset\preceq Q^{\circ}\prec P^{\circ}}{g}_{Q^{\circ}}(t^{2})\cdot(t^{2}-1)^{n-1-\dim(Q^{\circ})},\end{split}$
@@
This already recovers Stanley’s “master duality” theorem [29, Thm. 3.14.9], as well as the generalized Dehn-Sommerville relations for the $h$-polynomial of a lattice polytope (boiling down to Poincaré duality for $IH^{*}(X)$). In terms of the corresponding Ehrhart polynomials, this duality formula reads as
@@ 2. Homology Hirzebruch classes via mixed Hodge modules
commuting with proper pushforwards, with $Gr^{F}_{p}\hbox{\rm DR}({\mathcal{M}})\simeq 0$ for almost all $p$ and ${\mathcal{M}}$ fixed; see [23, Sec.2.3], [24, Sec.3.10, Prop.3.11] and [25, Sec.1]. Here $D^{b}_{\rm coh}(X)$ denotes the bounded derived category of ${\mathcal{O}}_{X}$-modules with coherent cohomology.
@@
$\begin{split}\hbox{\rm DR}_{y}([{\mathcal{M}}])&:=\sum_{i,p}\,(-1)^{i}\big{[}{\mathcal{H}}^{i}Gr^{F}_{-p}\hbox{\rm DR}({\mathcal{M}})\big{]}\cdot(-y)^{p}\\ &=\sum_{p}\,\big{[}Gr^{F}_{-p}\hbox{\rm DR}({\mathcal{M}})\big{]}\cdot(-y)^{p}\end{split}$
@@
(2.1) $\chi_{y}([H^{\bullet}]):=DR_{y}([H^{\bullet}])=T_{y*}([H^{\bullet}])=\sum_{j,p}\,(-1)^{j}\dim_{{\mathbb{C}}}\text{\rm Gr}_{F}^{p}H_{{\mathbb{C}}}^{j}\,(-y)^{p}$
@@
This follows, as in [19], from [17, Proposition 4.2], via the identification of these Ishida sheaves with the (shifted) graded parts of the filtered Du Bois complex (and similarly for the Du Bois complex of a complex algebraic $V$-manifold, by [21, Section 2.7]). ∎
@@
As shown in [27, Corollary 5.19, Remark 5.20], the transformations $\hbox{\rm DR}_{y}$ and $T_{y*}$ commute with duality, i.e.,
@@
and similarly for $T_{y*}$. A similar statement holds for the motivic transformations $mC_{y}$ and $T_{y*}$, with duality on the localization $K_{0}(var/X)[{\mathbb{L}}^{-1}]$ defined as in [27, Section 4B] and [3]. All these duality transformations are functorial for proper maps, so that for $X$ compact with $k:X\to pt$ proper, one gets
@@
In the case of a projective complex algebraic variety $X$, one has by the intersection cohomology Hodge index theorem [22, Sect.3.6] the equality
@@ 3. Computations via stratifications
Under the assumption of (virtual) stratum-wise constancy, formula (3.6) reduces the calculation of Hodge-Chern and Hirzebruch classes to the motivic calculus of [19], up to assigning the above mentioned Laurent polynomial weights $\chi_{y}(i_{s}^{*}[{\mathcal{M}}])$ to each stratum $S$, with $s\in S$ chosen. For instance, the stratum-wise constancy assumption is automatically satisfied in the toric context for ${\mathbb{T}}$-equivariant mixed Hodge modules, e.g., the $IC$-complex, see [30, Lemma 1.2].
@@
More generally, the property of stratum-wise constant higher direct image sheaves $R^{\ell}f_{*}{\mathbb{Q}}_{Y}|_{S}$ is satisfied if $f:f^{-1}(S)\to S$ is a Zariski locally trivial fibration for all strata $S\in{\mathcal{S}}$. This property also holds for a ${\mathbb{T}}$-equivariant proper morphism $f:Y\to X$, with $X$ a toric variety with torus ${\mathbb{T}}$ (by [30, Lemma 1.2]).
The terms $mC_{y}([S\hookrightarrow{\bar{S}}])=\hbox{\rm DR}_{y}\big{[}(i_{S,{\bar{S}}})_{!}{\mathbb{Q}}_{S}^{H}\big{]}$ appearing in formula (3.6) are computed via resolutions of singularities as follows (see [21, Prop.5.2.1] and [19, Prop.2.2]):
@@
(3.8) $mC_{y}([S\hookrightarrow{\bar{S}}])=\hbox{\rm DR}_{y}(\big{[}(i_{S,{\bar{S}}})_{!}{\mathbb{Q}}^{H}_{S}\big{]})=(\pi_{Z})_{*}\big{[}\mathcal{O}_{Z}(-D)\otimes\Lambda_{y}\Omega_{Z}^{1}(\log(D)\big{]}\in K_{0}({\bar{S}})[y],$
@@
For later use, let us also mention here the following formula, using the above notations (see, e.g., [27, Example 5.8]):
@@ 5. Characteristic class formulae in toric geometry
It suffices to prove the above properties on each open affine subvariety $U_{\sigma}\subset X$, for $\sigma\in\Sigma$, with the splitting $U_{\sigma}=Z_{\sigma}\times O_{\sigma},$ as recalled above. Let $p_{\sigma}:U_{\sigma}\to Z_{\sigma}$ be the smooth projection map with fiber $O_{\sigma}$. Then $(a)$ follows from the smooth base change property for the underlying constructible complex. Similarly, smooth pullback of mixed Hodge module complexes preserves the $K_{0}({\rm MHS}^{p})^{\rm Tate}$-module structure. Item $(b)$ can also be checked on the underlying perverse sheaf $IC_{X}$, where it follows by induction using $(a)$ and the Deligne construction. Again, these operations preserve the $K_{0}({\rm MHS}^{p})^{\rm Tate}$-module structure (see also [26, (1.7.6)]). For $(c)$, the stratum-wise constant case follows inductively from the fact that duality commutes (up to a shift) with the smooth pullback. Applying this to $[(j_{S})_{!}{\mathbb{Q}}^{H}_{X}]$ as in Example 3.5, one proves the virtual stratum-wise constant case since duality commutes with cross-products (and duality on a point preserves the Tate type). ∎
@@ 6. Applications to weighted lattice point counting
For a face $Q$ of $P$, denote by $k_{Q}:V_{\sigma_{Q}}:=X_{Q}\hookrightarrow X$ the inclusion of the orbit closure associated to the (cone $\sigma_{Q}$ of the) face $Q$. Note that we have $\dim_{\mathbb{R}}(Q)=\dim_{\mathbb{R}}(O_{\sigma_{Q}}).$ Then by (5.4) the following equality holds:
$\int_{X}ch({\mathcal{O}}_{X}(\ell D))\cap T_{y*}([{\mathcal{M}}])=\sum_{Q\preceq P}\chi_{y}(i_{x_{Q}}^{*}[{\mathcal{M}}])\cdot(1+y)^{\dim(Q)}\int_{X}ch({\mathcal{O}}_{X}(\ell D))\cap(k_{Q})_{*}td_{*}([\omega_{X_{Q}}]).$
@@
(6.7) $\int_{X}ch({\mathcal{O}}_{X}(\ell D))\cap(k_{Q})_{*}td_{*}([\omega_{X_{Q}}])=|{\rm Relint}(\ell Q)\cap M|.$
@@
(6.8) $\int_{X}ch({\mathcal{O}}_{X}(\ell D))\cap(k_{Q})_{*}td_{*}([\omega_{X_{Q}}])=\int_{X_{Q}}ch({\mathcal{O}}_{X_{Q}}(\ell D_{Q}))\cap td_{*}([\omega_{X_{Q}}]).$
The latter equality uses the projection formula, together with $(k_{Q})^{*}({\mathcal{O}}_{X_{P}}(\ell D_{P})={\mathcal{O}}_{X_{Q}}(\ell D_{Q})$. ∎
@@
Note that any ${\mathcal{M}}\in D^{b}{\rm MHM}(X_{P})$ (resp., $[{\mathcal{M}}]\in K_{0}({\rm MHM}(X_{P}))$), a (Grothendieck class of a) mixed Hodge module complex on $X_{P}$ which is (virtually) stratum-wise constant with respect to the stratification given by the torus orbits, induces a weight function on the faces of $P$ defined by $f_{Q}(y)=\chi_{y}(i_{x_{Q}}^{*}[{\mathcal{M}}])$, with $x_{Q}\in O_{\sigma_{Q}}$ a point in the orbit of $X_{P}$ associated to (the cone $\sigma_{Q}$ in the inner normal fan of $P$ corresponding to) the face $Q$ and $i_{x_{Q}}:\{x_{Q}\}\to X_{P}$ the inclusion map. Moreover, any weight function $f$ can be obtained in this way from such a mixed Hodge module complex ${\mathcal{M}}$, e.g., by using direct sums and shifts of $(j_{Q})_{!}{\mathbb{Q}}^{H}_{O_{\sigma_{Q}}}(n_{Q})$, with $n_{Q}\in{\mathbb{Z}}$ and $j_{Q}:O_{\sigma_{Q}}\hookrightarrow X$ the inclusion (or, using the subgroup $\chi_{Hdg}\left(\tau(F^{{\mathbb{T}}}(X)[y^{\pm 1}])\right)$ as in the Introduction). In particular, the expression $E_{P,f}(\ell,y)$ is obtained by evaluating a polynomial $E_{P,f}(z,y)$ at $z=\ell$.
@@
Let us next give some examples for Proposition 6.1, formulated in terms of Ehrhart theory. If $X=X_{P}$ is the projective toric variety with inner normal fan $\Sigma=\Sigma_{P}$ and ample Cartier divisor $D=D_{P}$ associated to a full-dimensional lattice polytope $P\subset M_{{\mathbb{R}}}\cong{\mathbb{R}}^{n}$, and $\ell\in{\mathbb{Z}}_{>0}$, for a face $Q$ of $P$, denote by $j_{Q}:O_{\sigma_{Q}}\hookrightarrow X$ the inclusion of the orbit associated to the (cone of the) face $Q$.
@@
    For ${\mathcal{M}}=(j_{Q})_{!}{\mathbb{Q}}^{H}_{O_{\sigma_{Q}}}$, we get
@@
    For ${\mathcal{M}}=(j_{Q})_{*}{\mathbb{Q}}^{H}_{O_{\sigma_{Q}}}$, we get by (5.2) and Danilov’s formula (6.1) that
@@
with $Q$ a face of $P$ and $Q^{\circ}$ the corresponding face of the polar polytope $P^{\circ}$, and $g_{Q^{\circ}}(t)\in{\mathbb{Z}}[t]$ the Stanley $g$-polynomial of $Q^{\circ}$, cf. [28, Sect.3]. We then get:
@@
(6.23) $\begin{split}\hbox{\rm DR}_{y}([{\mathcal{M}}])&=\sum_{Q\preceq P}\,\chi_{y}(i^{*}_{x_{Q}}[{\mathcal{M}}])\cdot(1+y)^{\dim(Q)}\cdot(k_{Q})_{*}[\omega_{V_{\sigma_{Q}}}]\\ &=\sum_{Q\preceq P}\,f_{Q}(y)\cdot\hbox{\rm DR}_{y}\left([(j_{Q})_{!}{\mathbb{Q}}^{H}_{O_{\sigma_{Q}}}]\right),\end{split}$
@@
$\hbox{\rm DR}_{y}\left({\mathbb{D}}_{X}\left([(j_{Q})_{!}{\mathbb{Q}}^{H}_{O_{\sigma_{Q}}}]\right)\right)=(-y)^{-\dim(Q)}\cdot\hbox{\rm DR}_{y}\left([(j_{Q})_{*}{\mathbb{Q}}^{H}_{O_{\sigma_{Q}}}]\right).$
@@
$\hbox{\rm DR}_{1/y}({\mathbb{D}}_{X}[{\mathcal{M}}])=\sum_{Q\preceq P}\,f_{Q}({y})\cdot(-y)^{\dim(Q)}\cdot\hbox{\rm DR}_{1/y}\left([(j_{Q})_{*}{\mathbb{Q}}^{H}_{O_{\sigma_{Q}}}]\right).$
@@
The reciprocity formula (6.20) can be applied to all cases of Example 6.5. In the (dual) cases $(a)$ and $(c)$, for ${\mathcal{M}}=(j_{Q})_{!}{\mathbb{Q}}^{H}_{O_{\sigma_{Q}}}$ with ${\mathbb{D}}_{X}[{\mathcal{M}}]=[(j_{Q})_{*}{\mathbb{Q}}^{H}_{O_{\sigma_{Q}}}(\dim(Q))]\in K_{0}({\rm MHM}(X))$, formula (6.20) specializes to the classical Ehrhart reciprocity (6.3) for $Q$. Indeed,